# Efficient AI in Practice: Training and Deployment of Efficient LLMs for Industry Applications

### v1

@@ Abstract.
Large language models (LLMs) have demonstrated remarkable performance across a wide range of industrial applications, from search and recommendations to generative tasks. Although scaling laws indicate that larger models generally yield better generalization and performance, their substantial computational requirements often render them impractical for many real-world scenarios at scale. In this paper, we present methods and insights for training small language models (SLMs) that deliver high performance and efficiency in deployment. We focus on two key techniques: (1) knowledge distillation and (2) model compression via quantization and pruning. These approaches enable SLMs to retain much of the quality of their larger counterparts while significantly reducing training, serving costs, and latency. We detail the impact of these techniques on a variety of use cases at a large professional social network platform and share deployment lessons—including hardware optimization strategies that enhance speed and throughput for both predictive and reasoning-based applications.
large language models, distillation, pruning, compression, quantization
1. Introduction
@@
Large language models (LLMs) (Dubey et al., 2024; Jiang et al., 2023; Team et al., 2023; Liu et al., 2024) have ushered in a new era in artificial intelligence and machine learning, driving significant improvements in deployed systems across various industries.
@@
LLMs suitable for real-world applications come in diverse forms, differing in size (ranging from hundreds of millions to hundreds of billions of parameters), architectural design (e.g., encoder-based models like BERT (Devlin, 2018) versus decoder-based models like GPT-3 (Brown et al., 2020)), and training paradigms (such as pre-training, instruction tuning, or test-time computation (Dubey et al., 2024; Team et al., 2023; Jiang et al., 2023; Mueller et al., 2023; Liu et al., 2024; Guo et al., 2025)).
In industry, LLMs are deployed to address a multitude of applications:
  • Search: Enabling embedding generation (Wang et al., 2022, 2023b) and semantic ranking and matching (Qin et al., 2023).
  • Retrieval and Ranking: Supporting retrieval (Zhao et al., 2024) and ranking of items (Li et al., 2023a; Firooz et al., 2025; Li et al., 2023b).
  • Generative Use Cases: Powering applications such as chatbots, assistants, and image generators (Achiam et al., 2023; Dam et al., 2024; Ramesh et al., 2022).
@@
Furthermore, scaling laws for LLMs have established a strong correlation between model size, validation loss, and downstream task performance (Kaplan et al., 2020; Hoffmann et al., 2022; Raffel et al., 2020; Wei et al., 2022). As a result, increasing the model size is often one of the most effective strategies for enhancing performance. Modern LLMs, particularly autoregressive decoder-only models, have expanded to hundreds of billions of parameters.
@@
While large LLMs exhibit high performance, deploying such large models incurs substantial infrastructure costs, especially for latency- or throughput-sensitive tasks. However, both academia and industry have developed strategies for creating and deploying efficient small language models (SLMs). Here, we primarily focus on methods which leverage an existing internally trained large LLM to create an efficient SLM that largely maintains the original model’s accuracy. Approaches to achieve this include white-box or black-box distillation (Hinton, 2015; Gu et al., 2024; Jin et al., 2021; Agarwal et al., 2024; Tunstall et al., 2023), compression techniques such as quantization (Frantar et al., 2022; Behdin et al., 2023) or sparsification (Frantar and Alistarh, 2023; Meng et al., 2024b; Sun et al., 2023; Meng et al., 2024a).
@@
In this work, we present a comprehensive set of lessons learned from developing and deploying various SLMs into production at a large-scale social networking company. We address a wide array of predictive and generative use cases, with inference performance and latency constraints in serving as key considerations. Specifically, our contributions are as follows:
@@
  • We discuss several large-scale use cases for which language models are useful.
@@
We apply these methods to the following real-world use cases, with experimental results and discussion provided in Section 4:
  • SLM for predictive tasks obtained through distillation and pruning - We leverage a foundation model (FM) for ranking and recommendations (Firooz et al., 2025; Li et al., 2023b) and leverage distillation and pruning to create an SLM that is efficient for serving latency-sensitive use cases. The final SLM we create is more than 20×20$\times$20 × smaller without any appreciable loss in quality.
  • SLM for reasoning task obtained through distillation - We leverage various flavors of KD to compress a latency-sensitive reasoning model by more than 5×5$\times$5 × without any appreciable loss in quality.
The remainder of the paper is organized as follows. In Section 2, we review related work. Section 3 details the efficiency techniques employed for creating, modifying, and deploying SLMs. Section 4 presents experimental results. We then address practical deployment considerations in Section 5 and conclude with a discussion of our findings and suggestions for future research.
2. Related work
Research on efficient LLMs has extensively explored pruning, knowledge distillation, and quantization. Muralidharan et al. (Muralidharan et al., 2024a) investigate pruning and distillation to maintain performance in smaller models. The Minitron paper (Muralidharan et al., 2024b) focuses on advanced pruning for deployment, while Apple’s study (Gunter et al., 2024) integrates multiple techniques, including quantization, for efficiency. Insights on hyper-parameter tuning (Ashkboos et al., 2024; Pareja et al., 2024) offer valuable training guidance but lack focus on compression or deployment, and surveys like (Zhu et al., 2023) provide overviews without empirical validation. In contrast, our work emphasizes lessons on efficiency through distillation and compression, coupled with practical deployment considerations such as inference optimization and GPU capacity utilization. With hundreds of millions of users worldwide, our use cases require models that efficiently handle large-scale data and deliver personalized, relevant content tailored to professional networking functionalities. We focus on fine-tuning language models specifically for industry use cases, ensuring alignment with our unique requirements and optimizing models for high-demand, real-world environments. This application-driven approach not only bridges the gap between theoretical compression methods and practical deployment needs but also provides actionable insights on optimizing GPU usage and deployment strategies. By addressing both efficiency and deployment challenges, our research distinguishes itself from more generalized studies, offering tangible solutions for large-scale social networks.
3. Methods
In this section, we detail various techniques that allow SLMs to retain strong generalization or task-specific performance, while allowing efficient serving from a latency or throughput standpoint. Specifically, we discuss training via knowledge distillation and post-training model compression. We also intersperse serving and training efficiency concerns throughout the write-up.
3.1. Knowledge Distillation
Knowledge distillation (KD) (Hinton, 2015) transfers knowledge from a larger and expressive “teacher” model to a smaller “student” model, allowing the latter to approximate teacher performance with reduced computational resources. KD can be broadly performed in two different ways (1) by leveraging the output of a teacher model to train the student (Tunstall et al., 2023; Guo et al., 2025)(also known as black-box distillation) or (2) by leveraging intermediate outputs (Muralidharan et al., 2024b; Hinton, 2015) (also known as white-box distillation). White-box distillation using the soft probabilistic outputs of the teacher is a powerful technique and helps provide richer information than hard labels used in supervised fine-tuning (SFT), helping the student generalize better, especially in tasks where smaller models struggle to discover patterns in noisy data.
(1) $\mathcal{L}[p_{\mathbf{y}},\mathcal{D}(p\|q_{\theta})]$ $=$
  1. (1)
    Responses are drawn from $p_{\mathbf{y}}$, which may correspond to ground truth data, the teacher model (sequence-level KD)  (Kim and Rush, 2016), or the student model itself (on-policy KD)  (Agarwal et al., 2024; Gu et al., 2024; Zhou et al., 2023). Recent advancements (Xu et al., 2024; Ko et al., 2024) explore a balance between on-policy and off-policy sampling to mitigate the mismatch between student-generated responses and the teacher’s distribution while addressing inefficiencies in online student autoregressive training.
  2. (2)
  • Forward Kullback-Leibler (KL) Divergence (FKL):
    $\mathcal{D}_{\text{FKL}}\left[p(y_{t}|\mathbf{y}_{<t},\mathbf{x})\|q_{\theta}(y_{t}|\mathbf{y}_{<t},\mathbf{x})\right]=\sum_{i\in\mathcal{V}}p(i|\cdot)\log\left(\frac{p(i|\cdot)}{q_{\theta}(i|\cdot)}\right),$
  • Reverse Kullback-Leibler Divergence (RKL):
    $\mathcal{D}_{\text{RKL}}\left[p(y_{t}|\mathbf{y}_{<t},\mathbf{x})\|q_{\theta}(y_{t}|\mathbf{y}_{<t},\mathbf{x})\right]=\sum_{i\in\mathcal{V}}q_{\theta}(i|\cdot)\log\left(\frac{q_{\theta}(i|\cdot)}{p(i|\cdot)}\right),$
  • Jensen-Shannon Divergence (JSD):
    $\mathcal{D}_{JS(\beta)}\left[p(y_{t}|\mathbf{y}_{<t},\mathbf{x})\|q_{\theta}(y_{t}|\mathbf{y}_{<t},\mathbf{x})\right]=\;$ $\beta\mathcal{D}_{\text{FKL}}\left[p\|m\right]$
    $+(1-\beta)\mathcal{D}_{\text{FKL}}\left[q_{\theta}\|m\right],$
    where $m=\beta p+(1-\beta)q_{\theta}$ is the mixture distribution.
3.2. Post-training compression
Model compression is a widely-studied area of machine learning efficiency. We specifically focus on post-training compression (PTC) techniques to improve the inference efficiency of SLMs and LLMs.
Our focus is on post-training compression, where compression is applied to the model after training. Recent compression methods follow a layerwise compression approach (Equation 2) since modern models have ballooned in size and it can be scaled to large models without a major compromise in accuracy.
We now present a mathematical framework for layerwise PTC using calibration data. Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ denote the calibration data that serve as inputs to a linear layer of the model (e.g., an MLP or an attention projection). Here, $n$ is the number of tokens in the calibration dataset, and $d$ is the input dimension of the layer. For instance, in the case of an MLP down projection layer of a Transformer block, $d$ corresponds to the intermediate size of the model.
(2) $\min_{\hat{\mathbf{W}}}\|\mathbf{X}\mathbf{W}-\mathbf{X}\hat{\mathbf{W}}\|_{F}^{2}\quad\text{subject to}\quad\hat{\mathbf{W}}\in\mathcal{Q},$
where $\mathcal{Q}\subseteq\mathbb{R}^{d\times p}$ denotes the set of feasible solutions that conform to a particular compression scheme. In practice, the set $\mathcal{Q}$ often exhibits a discrete structure, which renders the optimization problem in (2) challenging to solve.
Here, we consider two compression techniques:
Quantization In quantization, the model weights are represented in lower precision, using a fewer number of bits. Quantization has proved to be successful in the LLM domain, to obtain compressed models with small accuracy loss (Frantar et al., 2022; Xiao et al., 2023). In this work, we consider weight-only quantization, where only model weights are quantized, as well as weight and activation quantization. We study methods such as GPTQ (Frantar et al., 2022) and QuantEase (Behdin et al., 2023) for 4-bit weight-only quantization (aka W4A16), SmoothQuant (Xiao et al., 2023) for 8-bit weight-and-activation quantization (aka W8A8), and 8-bit floating-point (FP8) quantization. Since quantization is dependent on hardware, we discuss the details of quantization-related experiments in Section 5 (deployment).
Structured Pruning Neural network pruning (LeCun et al., 1989; Hassibi et al., 1993) is a post-training compression technique, where some model weights are set to zero, resulting in sparse model weight matrices. Recent years have seen a surge of papers in pruning techniques for LLMs—examples include (Meng et al., 2024a; Frantar and Alistarh, 2023; Sun et al., 2023). Without any structure on the model sparsity, however, it can be difficult to realize any inference acceleration from pruning. Therefore, in this work, we pursue a structured pruning approach. In structured pruning, the goal is to obtain smaller models via removing some neurons from the model weights. In particular, we study MLP pruning, where the goal is to reduce the intermediate size of the model via removing some hidden neurons in feed-forward layers. We also study attention pruning, where we remove a certain number of attention heads from the model (Meng et al., 2024b; Kwon et al., 2022; Kurtić et al., 2024). In this paper, we use OSSCAR (Meng et al., 2024b) which uses a discrete optimization approach for structured pruning. OSSCAR can be scaled to large scale problems, which we consider. We leverage OSSCAR because it represents the state-of-the-art when it comes to post-training structured pruning of LLMs and leads to the least drop in accuracy when compared to other methods.
Figure 1. Overview of the process of creating SLMs via distillation and compression.
3.3. Training and serving efficiency
Training Efficiency LLM training presents a formidable challenge due to the sheer scale of these models and the quadratic complexity of transformer architectures. Model FLOPs utilization (MFU) (Chowdhery et al., 2023) is commonly used to measure GPU efficiency, making it necessary to optimize kernel operations and distributed training strategies. We have implemented Liger Kernel (Hsu et al., 2024) in Triton (Tillet et al., 2019), incorporating several key optimizations. First, we employ kernel fusion to reduce repetitive memory transfers between SRAM and DRAM. Next, we adopt in-place tensor modifications to avoid creating additional tensors whenever possible, thus lowering the memory footprint. We also apply chunking, which prevents the full materialization of large tensors and provides tuning flexibility while maintaining comparable performance. In combination, these optimizations reduce training time by 20% and memory usage by 60%. Additional performance-memory tradeoffs such as gradient checkpointing and CPU offloading can lead to as much as a threefold speedup.
For distributed training, we use ZeRO (Rajbhandari et al., 2020) to shard model parameters and data across multiple GPUs, overlapping computation with communication to sustain high MFU. Together with the DeepSpeed team, we have optimized the ZeRO algorithm, and for network-constrained clusters, we have developed ZeRO++ (Wang et al., 2023a; Dai et al., 2024) to mitigate non-deterministic synchronization issues that can hinder convergence. ZeRO++ provides a 2.4$\times$ speedup over vanilla ZeRO.
Serving Efficiency Serving LLMs efficiently poses unique challenges due to both high computational demands and strict latency requirements. In production environments, the choice of a serving framework is pivotal for maximizing throughput and minimizing response times. Several solutions, including vLLM, SGLag, TRT-LLM, and MLC-LLM (Kwon et al., 2023; Zheng et al., 2024; NVIDIA, 2024; team, 2025), have been proposed to address these needs. In our use cases, we evaluated vLLM and SGLang extensively. Our benchmarks revealed that SGLang is better suited to our workloads because its radix tree-based caching mechanism aligns well with our usage patterns and it integrates tightly with FlashInfer (Ye et al., 2025), whose efficient attention kernels accelerate the sequence lengths and batch sizes we typically handle.
To further improve serving performance, we deploy our models on NVIDIA H100 GPUs at FP8 precision, striking a practical balance between computational efficiency and model quality. Additional details regarding the serving engine configurations for our experiments can be found in Section 5.
4. Experiments
4.1. Setup
@@
Prompt Structures For predictive tasks, we are mainly interested in ranking use cases. Hence, the use of decoder-based LLMs is prefill-dominant. For generative and reasoning based-tasks, we are also interested in decoding latencies. We relegate the details of the prompt structures to subsections discussing the specific use cases.
@@ 4.2.1. Problem setup
We base our experiments on an internal foundation model (FM) trained using text, primarily for the purpose of ranking and recommendations  (Firooz et al., 2025). The FM consists of 100B+ parameters and is trained to approximate the following distribution for a large variety of recommendation tasks where users interact with items:
(3) $P(m,(e_{1},t_{1}),...,(e_{T},t_{T})),$
@@
where $m$ represents a user, and each pair $(e_{t},i_{t})$ for $t=1,...,T$ represents the user’s interaction $i_{t}$ (like or click or equivalent) with an entity $e_{t}$ (such as post on a social media platform) . As mentioned before, the featurization is performed purely via text, allowing the FM to effectively generalize across heterogeneous tasks. The model is then used to estimate the following probabilities for future interactions with entities:
(4) $P(i_{t},i_{t+1},...|\text{Task instruction},m,(e_{1},i_{1}),...,(e_{t-1},i_{t-1}),e_{t},e_{t+1},...)$
@@
Text-based featurization makes decoder-only LLMs an attractive option for training this model jointly on a large and varied of recommendation tasks. The FM uses single-token generation for pointwise ranking and probability estimation. The reader is encouraged to peruse  (Firooz et al., 2025) for more details about the FM.
Due to the large size of the FM—which contains over 100 billion parameters—serving it online for latency-sensitive applications is challenging. In this work, we present experiments demonstrating how we achieved a more than 20$\times$ reduction in model size, enabling the online serving of a distilled version of the FM with only a modest loss in accuracy.
@@
Our approach, illustrated in Figure 1, proceeds in three stages: 1) Distillation on the full model. 2) One-shot pruning111“One-shot” indicates pruning without any re-training. to significantly reduce the model size. 3) Re-distillation of the pruned model to recover generalization capabilities. All distillation is performed in a post-training setting using data from various recommendation tasks. We optionally also quantize the model (details in Section 5). We measure quality by reporting the AUC on the test sets of in-domain tasks, computing AUC per task and then averaging across tasks.
4.2.2. Distillation Results
To evaluate how knowledge distillation (KD) affects model generalization, we compare it against standard supervised fine-tuning (SFT). We focus on the task performance retention of the distilled (or fine-tuned) student models relative to the original foundation model (FM). Specifically, we use Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct (Dubey et al., 2024) as our student models. These models offer strong performance while remaining sufficiently compact for throughput- and latency-sensitive environments. In both KD and SFT, responses are generated from the ground-truth action-prediction labels.
For KD, the per-token loss is a weighted combination of: 90% forward KL divergence between teacher and student logits, and 10% cross-entropy loss with the ground-truth labels.
Additionally, we include an extra 5% loss contribution computed over the entire sequence (including the prompt), normalized by its token count. Thus, 95% of the loss is computed solely on action prediction tokens, while the remaining 5% is computed over the prompt tokens.
Each KD and SFT configuration undergoes hyperparameter tuning (e.g., peak learning rate, warmup schedule, decay schedule, and weight decay). Figure 2 summarizes the results by reporting the AUC delta relative to the original FM (which achieves the best performance). The main observations are as follows:
  • SFT: The 3B and 8B student models fine-tuned only with SFT underperform compared to the FM, which is expected given their smaller size and post-training. The performance drop for the 3B model (-1.21%) is larger than that of the 8B model (-0.62%).
  • KD: Using logit supervision from the FM consistently preserves task performance better than SFT. The 8B-KD model shows a minor AUC drop of -0.06% (compared to -0.62% for 8B-SFT), while the 3B-KD model (-0.15%) substantially mitigates the loss relative to 3B-SFT (-1.21%). These results demonstrate the effectiveness of KD for transferring knowledge, allowing smaller models to remain competitive.
Overall, KD performs well on transferring the knowledge from the teacher model and improving smaller model’s generalization performance, enabling us to compress a 100B+ FM into smaller models with 3B and 8B parameters.
8B-KD8B-SFT3B-KD3B-SFT11-1- 10.50.5-0.5- 0.5000.060.06-0.06- 0.060.620.62-0.62- 0.620.150.15-0.15- 0.151.211.21-1.21- 1.21AUC Delta (%)
Figure 2. Comparison of Distillation and SFT on the Foundation Model. Knowledge distillation consistently outperforms SFT by effectively leveraging teacher supervision to preserve and enhance performance.
4.2.3. Post-training Compression Results
After obtaining the distilled student models, we apply structured pruning and on-the-fly FP8 quantization to further compress the model to meet our serving latency requirements (steps outlined in Figure 1). Since the distilled models (DMs) are decoder-only transformers, our approach focuses on applying structured pruning to remove redundant MLP up/down projection neurons, as well as attention heads in each transformer layer while preserving its capabilities on in-domain ranking tasks. Since one-shot pruning results in loss in quality, we apply targeted fine-tuning after each pruning step to ensure that the model adapts to its reduced size without significant performance loss. To this end, we again leverage knowledge distillation to bridge the gap between the original and pruned models, transferring key insights and ensuring the pruned version closely aligns with the outputs of the original foundational model.
For structured pruning, we leverage the OSSCAR algorithm (Meng et al., 2024b). We now discuss the details of various ablations that we conducted towards structured pruning.
We also explore gradual pruning (Benbaki et al., 2023) in which the model is pruned in multiple steps, with knowledge distillation after each pruning. Table 3 summarizes results for pruning a distilled 3B model to 2.4B (a 37.5% MLP sparsity) either in a single step or in two smaller steps (removing 1536 hidden neurons in each step, for a total of 3072). The gradual approach recovers model AUC better than the single-step approach, achieving near-lossless compression from 3B to 2.4B.
Model AUC Delta (%)
8B Distilled Model -
6.4B Pruned Model (20%) + SFT -0.47%
6.4B Pruned Model (20%) + Distillation -0.06%
Table 1. Distillation vs. SFT for post-pruning retraining. The pruned model has its MLP layers pruned by $20\%$.
Model AUC Delta (%)
8B Model -
6.8B Pruned Model 0.0%
6.4B Pruned Model -1.33%
6.0B Pruned Model -1.72%
Table 2. Evaluation of the 8B-parameter model post-SFT and its pruned variants, focusing on MLP pruning performed in a one-shot manner (i.e., no retraining after pruning).
Model #Params AUC Delta (%)
3B (Distilled from FM) 3B -
MLP Prune + Distill 2.4B -0.12%
MLP Prune + Distill (Gradual) 2.4B 0.03%
Table 3. Comparison of one-step vs gradual pruning for a 3B model distilled from the FM.
Model #Params AUC Delta (%)
MLP Pruning 2.4B -
$\wedge$ + Attention Pruning 2.1B -1.07%
$\wedge$ + Distillation 2.1B 0.02%
Table 4. Results for attention pruning. We consider the 2.4B gradual pruning model from Table 3 as the base. The second row shows the result for one-shot attention pruning, while the last row shows the results after performing distillation.
Full PrecisionC4 1024C4 4096Task-Specific 350Task-Specific 70011-1- 10.80.8-0.8- 0.80.60.6-0.6- 0.60.40.4-0.4- 0.40.20.2-0.2- 0.200001.061.06-1.06- 1.060.640.64-0.64- 0.640.120.12-0.12- $-0.1$AUC Delta (%)
Figure 3. Comparison of one-shot pruning methods. The bars indicate the drop (in percentage points) relative to the full precision baseline. The pruned model is a 6.4B model (20% MLP pruning).
Effect of calibration data Figure 3 captures the effect of the calibration dataset $\boldsymbol{X}$ (discussed in Section 3 and Equation (2)) on the accuracy of the pruned versions of the 8B student model (results for the 3B model are similar and are hence omitted for clarity). The Full Precision bar illustrates the baseline accuracy of the non-pruned model. We consider two different datasets for calibration - C4 (Raffel et al., 2020), an open source dataset, and an in-domain dataset. When we prune using a randomly sampled portion of the C4 dataset (1,024 or 4,096 examples), accuracy drops, although more samples mitigate this drop to an extent. These results indicate that increasing the number of calibration examples from 1,024 to 4,096 can partially recover lost accuracy due to pruning. However, leveraging fewer but more domain-specific samples (350 or 700 examples from the target task) yields better accuracy values, which closely match the full precision baseline. This highlights the importance of using task-relevant data for calibration, even if it involves fewer examples, as it can produce more accurately pruned models than generic calibration sets.
4.3. Reasoning Task

### v2

# Scaling Down, Serving Fast: Compressing and Deploying Efficient LLMs for Recommendation Systems

@@ Abstract
Large language models (LLMs) have demonstrated remarkable performance across a wide range of industrial applications, from search and recommendation systems to generative tasks. Although scaling laws indicate that larger models generally yield better generalization and performance, their substantial computational requirements often render them impractical for many real-world scenarios at scale. In this paper, we present a comprehensive set of insights for training and deploying small language models (SLMs) that deliver high performance for a variety of industry use cases. We focus on two key techniques: (1) knowledge distillation and (2) model compression via structured pruning and quantization. These approaches enable SLMs to retain much of the quality of their larger counterparts while significantly reducing training/serving costs and latency. We detail the impact of these techniques on a variety of use cases in a large professional social network platform and share deployment lessons, including hardware optimization strategies that improve speed and throughput for both predictive and reasoning-based applications in Recommendation Systems.
Scaling Down, Serving Fast: Compressing and Deploying Efficient LLMs for Recommendation Systems
Kayhan Behdin1,3,5, Ata Fatahibaarzi 1,3,5, Qingquan Song2,3,5, Yun Dai 2,3,5, Aman Gupta 2,3,5, Zhipeng Wang ,1,3, Shao Tang 1,3, Hejian Sang 1,3, Gregory Dexter 1, Sirou Zhu 1, Siyu Zhu1, Tejas Dharamsi1, Vignesh Kothapalli2, Zhoutong Fu1, Yihan Cao1, Pin-Lun Hsu2, Fedor Borisyuk1, Natesh Pillai1, Luke Simon2, Rahul Mazumder 1,3,4 1 LinkedIn, 2 Work done at LinkedIn, 3 LinkedIn LLM Efficiency Core Team, 4 MIT 5 Authors contributed equally to this research, $\ast$ Correspondence: zhipwang@linkedin.com
1 Introduction
@@
Large language models (LLMs) Dubey et al. (2024); Jiang et al. (2023); Team et al. (2023); Liu et al. (2024) have ushered in a new era in artificial intelligence and machine learning, driving significant improvements in deployed systems across various industries.
@@
LLMs suitable for real-world applications come in diverse forms, differing in size (ranging from hundreds of millions to hundreds of billions of parameters), architectural design (e.g., encoder-based models like BERT Devlin (2018) versus decoder-based models like GPT-3 Brown et al. (2020)), and training paradigms (such as pre-training, instruction tuning, or test-time computation Dubey et al. (2024); Team et al. (2023); Jiang et al. (2023); Mueller et al. (2023); Liu et al. (2024); Guo et al. (2025)).
Specifically for the Social Network Platforms, LLMs are heavily leveraged and deployed for a multitude of applications: 1. Semantic Search (e.g., embedding generation Wang et al. (2022, 2023b) and semantic ranking/matching Qin et al. (2023)); 2. Recommendation Systems (RecSys), specifically Retrieval and Ranking Zhao et al. (2024) Li et al. (2023b); Firooz et al. (2025); Li et al. (2023a); 3. Generative Use cases, such as chatbots, assistants, image generators, etc. Achiam et al. (2023); Dam et al. (2024); Ramesh et al. (2022).
@@
Furthermore, scaling laws for LLMs have established a strong correlation between model size, validation loss, and downstream task performance Kaplan et al. (2020); Hoffmann et al. (2022); Raffel et al. (2020); Wei et al. (2022). As a result, increasing the size of the model is often one of the most effective strategies to enhance performance. Modern LLMs, particularly autoregressive decoder-only models, have expanded to hundreds of billions of parameters.
@@
Although large LLMs exhibit extraordinary performance, the deployment of such large models incurs substantial infrastructure costs, especially for latency- or throughput-sensitive tasks in Recommendation Systems (RecSys). However, both academia and industry have developed strategies for creating and deploying efficient small language models (SLMs). Here, we primarily focus on methods that leverage an existing internally trained large LLM to create an efficient SLM that largely maintains the original model’s accuracy. Approaches to achieve this include white-box or black-box distillation Hinton (2015); Gu et al. (2024); Jin et al. (2021); Agarwal et al. (2024); Tunstall et al. (2023), compression techniques such as quantization Frantar et al. (2022); Behdin et al. (2023) and sparsification Frantar and Alistarh (2023); Meng et al. ; Sun et al. (2023); Meng et al. (2024a); Campos et al. (2023).
@@
In this work, we present a suite of insights from the training and deployment of various efficient SLMs in production at a large-scale professional social networking company. We address a wide array of predictive and generative use cases in RecSys (including ranking, recommendations, and reasoning), with inference performance and latency constraints in serving as key considerations. Our contributions are as follows.
@@
  • We discuss several large-scale RecSys use cases for which language models are useful.
@@
2 Preliminaries
@@
Prompt Structures For predictive tasks, we are mainly interested in ranking use cases. Hence, the use of decoder-based LLMs is prefill-dominant. For generative and reasoning based-tasks, we are also interested in decoding latencies. We relegate the details of the prompt structures to subsections for the specific use cases.
@@
Foundational model for RecSys We base our experiments on an internal foundation model (FM) trained using text, primarily for the purpose of ranking and recommendations  Firooz et al. (2025). The FM is a Mixture-of-Experts (MoE) model with an architecture motivated by the Mixtral family (Jiang et al., 2024). In particular, each expert is initialized based on a Llama 3.1 8B Instruct model (Dubey et al., 2024), with 16 experts in total (4 active per token). The FM is trained to approximate the following distribution for a large variety of recommendation tasks where users interact with items:
$P(m,(e_{1},t_{1}),...,(e_{T},t_{T})),$ (1)
@@
where $m$ represents a user, and each pair $(e_{t},i_{t})$ for t=1,,Tt=1,...,T represents the user’s interaction $i_{t}$ (like or click or equivalent) with an entity $e_{t}$ (such as post on a social media platform) . As mentioned before, the featurization is performed purely via text, allowing the FM to effectively generalize across heterogeneous tasks. The model is then used to estimate the following probabilities for future interactions with entities:
$P(i_{t},i_{t+1},\dots\mid\text{Task instruction},m,\\ (e_{1},i_{1}),\dots,(e_{t-1},i_{t-1}),e_{t},e_{t+1},\dots)$ (2)
@@
Text-based featurization makes decoder-only LLMs an attractive option for training this model jointly on a large and varied of recommendation tasks. The FM uses single-token generation for pointwise ranking and probability estimation. The reader is encouraged to peruse the paper Firooz et al. (2025) for more details about the FM.
[…]
Figure 1: Overview of the process of creating SLMs via distillation and compression.
3 Methodology
We apply these methods to the following real-world use cases in RecSys, with detailed methodology described in Appendix A.
  • SLM for predictive tasks obtained through distillation and pruning - We leverage a foundation model (FM) for ranking and recommendations Firooz et al. (2025); Li et al. (2023a) and leverage distillation and pruning to create an SLM that is efficient for serving latency-sensitive use cases. The final SLM we create is more than $20\times$ smaller without any appreciable loss in quality.
  • SLM for reasoning task obtained through distillation - We leverage various flavors of KD to compress a latency-sensitive reasoning model by more than $5\times$ with comparable quality.
@@ 3 Methodology
Our approach, illustrated in Figure 1, proceeds in three stages: 1) Distillation on the full model. 2) One-shot structured pruning111“One-shot” indicates pruning without re-training. to significantly reduce the model size. 3) Re-distillation of the pruned model to recover generalization capabilities. All compression is performed in a post-training setting using data from various recommendation tasks (see Section A for details on methods). We optionally also quantize the model (details in Section 6). We measure quality by reporting the AUC on the test sets of in-domain tasks, computing AUC per task and then averaging across tasks.
4 Experimental Results for Reasoning Tasks